# A Unified Implicit Attention Formulation for Gated-Linear Recurrent Sequence Models

### v1

@@ Appendix B Weakly Supervised Semantic Segmentation
In weakly supervised semantic segmentation (WSSS), a common approach involves first training a classifier on image-level labels and then extracting Class Activation Maps (CAMs) for individual images, which highlight regions that the classifier deems relevant to specific classes. The state-of-the-art methods then employ these CAMs as pseudo-masks to train a segmentation decoder.
In this context, we adopt our proposed Mamba-Attr XAI method for vision-mamba models. We assess its competitiveness against the well-established CAMs in generating pseudo-labels for transformers. To ensure a fair comparison, we fine-tune both DeiT-Small and Vision Mamba-Small models under identical conditions over the Pascal-voc 2012 [17] dataset, excluding multi-scale training, inference, or any other modifications. This controlled setting isolates the influence of our Mamba-Attr method on the quality of the generated pseudo-labels.
The results are presented in Table 4. Evidently, Mamba-Attr XAI method achieves competitive results surpassing the baseline approach of Class Activation Maps (CAMs) without any additional modifications. This is evident in the mean Intersection-over-Union (mIoU) score, where Mamba-Attr ($52.11\%$) outperforms the CAM of DeiT-Small ($35.99\%$) by a sizable gap. While Mamba-Attr does not reach the state-of-the-art performance of Toco [49] ($61.10\%$), it achieves, out of the box, a substantial improvement over CAM and comes surprisingly close to this much more elaborate multi-phase learning method which utilizes multiple loss terms specifically designed to enhance the quality of the initial CAM map. These results suggest that Mamba-Attr XAI offers a powerful and efficient solution for WSSS tasks with vision-mamba models.
Table 4: Evaluation and comparison of the pseudo-labels for the different classes in Pascal-voc 2012 [17]. Results are in mIoU
@@ Appendix C Perturbation Experiments for NLP
In Figure 6, we present perturbation results for both positive and negative settings. We utilize Mamba [^1] https://huggingface.co/trinhxuankhai/mamba_text_classification, RWKV [^2] https://huggingface.co/BlinkDL/rwkv-2-pile-430m and BERT [^3] https://github.com/hila-chefer/Transformer-Explainability models as the models of interest and fine-tune them on the IMDB sentiment classification dataset, employing the [CLS] token for all samples. Subsequently, we evaluate the explanation quality using a similar procedure as proposed in [2] for the perturbation experiment.
The results reveal that Mamba-attr, based on our new attention formulation, achieves superior AUC for both negative and positive perturbations compared to the previous attention formulation by [2]. Additionally, our unified attention formulation is effective for RWKV models, yielding comparable results to those of Mamba and BERT.
Moreover, as an ablation, the first column of Fig. 6 demonstrates that including the gate branch, as presented in our full method, consistently improves performance.
@@ Appendix D Additional qualitative results for NLP
Additional NLP results obtained on IMDB dataset are presented in Figure 5. In panel (a), we show the results for the previously proposed Mamba’s attention [2]. Panel (b) shows our proposed Mamba’s attention. Lastly, panel(c) presents our proposed method over RWKV. In red, we show a negative sentiment, and in blue, we show a positive sentiment.
As can be seen from these qualitative results, the explanation maps generated by our new attention formulation exhibit sparser and more accurate heatmaps of relevant words than those of Ali et al. [2], aligning with the desired properties of XAI methods. Similarly, the results for RWKV models show comparable success to those of Mamba.
@@
Table 5: Additional qualitative results on the IMDB dataset. (a) The Mamba attention of Ali et al. [2]. (b) Our Mamba attention method. (c) Our RWKV attention.

### v2

@@ Appendix B Weakly Supervised Semantic Segmentation
In weakly supervised semantic segmentation (WSSS), a common approach involves first training a classifier on image-level labels and then extracting Class Activation Maps (CAMs) for individual images, which highlight regions that the classifier deems relevant to specific classes. The SoTA methods then employ these CAMs as pseudo-masks to train a segmentation decoder.
In this context, we adopt our proposed Mamba-Attr XAI method for vision-Mamba models. We assess its competitiveness against the well-established CAMs in generating pseudo-labels for Transformers. To ensure a fair comparison, we fine-tune both DeiT-Small and ViM-Small models under identical conditions over the Pascal-voc 2012 (Everingham et al., 2010) dataset, excluding multi-scale training, inference, or any other modifications. This controlled setting isolates the influence of our Mamba-Attr method on the quality of the generated pseudo-labels.
The results are presented in Table 5. Evidently, Mamba-Attr XAI method achieves competitive results surpassing the baseline approach of Class Activation Maps (CAMs) without any additional modifications. This is evident in the mean Intersection-over-Union (mIoU) score, where Mamba-Attr ($52.11\%$) outperforms the CAM of DeiT-Small ($35.99\%$) by a sizable gap. While Mamba-Attr does not reach the state-of-the-art performance of Toco (Ru et al., 2023) ($61.10\%$), it achieves, out of the box, a substantial improvement over CAM and comes surprisingly close to this much more elaborate multi-phase learning method which utilizes multiple loss terms specifically designed to enhance the quality of the initial CAM map. These results suggest that Mamba-Attr XAI offers a powerful and efficient solution for WSSS tasks with vision-Mamba models.
Table 5: Evaluation and comparison of the pseudo-labels for the different classes in Pascal-voc 2012 (Everingham et al., 2010). Results are in mIoU
@@ Appendix C Perturbation Experiments for NLP
In this section, we present results for the perturbation test in the NLP domain with fine-tuned classifiers. In this setting, we fine-tune the last layers of various LLMs and append the [CLS] token to all samples to generate explanation maps, similar to the methods used in vision models.
In Figure 5, we present perturbation results for both positive and negative settings. We utilize Mamba [^1] https://huggingface.co/trinhxuankhai/mamba_text_classification, RWKV [^2] https://huggingface.co/BlinkDL/rwkv-2-pile-430m and BERT [^3] https://github.com/hila-chefer/Transformer-Explainability models as the models of interest and fine-tune them on the IMDB sentiment classification dataset, employing the [CLS] token for all samples. Subsequently, we evaluate the explanation quality using a similar procedure as proposed in (Ali et al., 2024) for the perturbation experiment.
The results reveal that Mamba-attr, based on our new attention formulation, achieves superior AUC for both negative and positive perturbations compared to the previous attention formulation by (Ali et al., 2024). Additionally, our unified attention formulation is effective for RWKV models, yielding comparable results to those of Mamba and BERT.
Moreover, as an ablation, the first column of Figure 5 demonstrates that including the gate branch, as presented in our full method, consistently improves performance.
@@ Appendix D Additional qualitative results for NLP
Additional NLP results obtained on IMDB dataset are presented in Figure 6. In panel (a), we show the results for the previously proposed Mamba’s attention (Ali et al., 2024). Panel (b) shows our proposed Mamba’s attention. Lastly, panel(c) presents our proposed method over RWKV. In red, we show a negative sentiment, and in blue, we show a positive sentiment.
As can be seen from these qualitative results, the explanation maps generated by our new attention formulation exhibit sparser and more accurate heatmaps of relevant words than those of Ali et al. (2024), aligning with the desired properties of XAI methods. Similarly, the results for RWKV models show comparable success to those of Mamba.
@@
Table 6: Additional qualitative results on the IMDB dataset. (a) The Mamba attention of Ali et al. (2024). (b) Our Mamba attention method. (c) Our RWKV attention.